# Steering Towards Fairness: Mitigating Political Bias in LLMs

### v2

@@ 2.2 Steering Vectors and Ensemble Approaches for Mitigation
To address these limitations, our work introduces an activation-based bias mitigation pipeline tailored for decoder LLMs. Our core contributions include: A scalable, multilingual framework for layer-wise activation extraction using PCT-based contrastive pairs; A representation-level analysis method that identifies and aggregates ideological bias directions. And the first integration of Steering Vector Ensembles into decoder-based LLMs for mitigating political bias across both social and economic axes. By bridging output-based evaluation and internal mitigation strategies, we provide a new foundation for probing and correcting ideological bias in multilingual generative models. Although we focus on Political Compass Test (PCT) prompts, our pipeline is modular and could be extended to other domains such as healthcare or education. The ensemble design also improves robustness to prompt framing by aggregating across multiple paraphrases.
@@ 3.3 Target Model and Layer Selection
We selected deepseek-llm-7b-chat [^1] https://huggingface.co/deepseek-ai/deepseek-vl-7b-chat, Mistral model due to its strong multilingual capabilities and transparent architecture. We also evaluate Mistral-7B-v0.1222https://huggingface.co/mistralai/Mistral-7B-v0.1 to compare bias behavior across model families. We selected layers 8, 12, 16, 20, 24, as layer-wise profiling indicated that mid-level layers encode the strongest ideological signals, whereas early layers capture lexical patterns and very late layers primarily influence fluency.
@@ 5.1 High Resource Language
Figure 7: Bias reduction performance of SVE and ISV across different contrastive pair types.

### v3

@@ 2.2 Steering Vectors and Ensemble Approaches for Mitigation
To address these limitations, our work introduces an activation-based bias mitigation pipeline tailored for decoder LLMs. Our core contributions include: A scalable, multilingual framework for layer-wise activation extraction using PCT-based contrastive pairs; A representation-level analysis method that identifies and aggregates ideological bias directions. And the first integration of Steering Vector Ensembles into decoder-based LLMs for mitigating political bias across both social and economic axes. By bridging output-based evaluation and internal mitigation strategies, we provide a new foundation for probing and correcting ideological bias in multilingual generative models. Although we focus on Political Compass Test (PCT) prompts, our pipeline is modular and could be extended to other domains such as healthcare or education. The ensemble design also improves robustness to prompt framing by aggregating across multiple paraphrases. Our implementation is publicly available to support reproducibility and further work  [^1] https://github.com/Afx-Msh/SVE_Mitigation.
@@ 3.3 Target Model and Layer Selection
We selected deepseek-llm-7b-chat [^2] https://huggingface.co/deepseek-ai/deepseek-vl-7b-chat, Mistral model due to its strong multilingual capabilities and transparent architecture. We also evaluate Mistral-7B-v0.1333https://huggingface.co/mistralai/Mistral-7B-v0.1 to compare bias behavior across model families. We selected layers 8, 12, 16, 20, 24, as layer-wise profiling indicated that mid-level layers encode the strongest ideological signals, whereas early layers capture lexical patterns and very late layers primarily influence fluency.
@@ 5.1 High Resource Language
Figure 7: Bias reduction performance on Deep Seek model for SVE and ISV across different contrastive pair types.
@@
The results show in the Figure  8 that both the Urdu and Punjabi models handle bias reduction well while keeping their answers natural. They cut out biased keywords to a moderate degree about 0.6 on the scale without over filtering. Response quality stays high, around 0.85 - 0.9, and the overall flow of the replies (coherence) remains steady for both languages. The scatter plot on the right makes it clear that when the overall debiasing score goes up, the quality of the responses also rises, meaning stronger bias mitigation doesn’t hurt the readability or sense of the output.
Figure 8: Performance of Bias-Mitigation Methods for Urdu and Punjabi. The left plot compares Keyword Reduction, Response, Quality, and coherence for ISV and SVE.